OpenStudy (anonymous):

Will fan and medal! What is the probability of getting 4 heads when tossing a coin 6 times?

OpenStudy (anonymous):

tttt, tftf, ffff, ftff, tfff, tttf so you have 25% cause its a 50/50 but four times in a row takes out a 25 so i would say 1/6 change of doing that

OpenStudy (anonymous):

i hoped i help haha

OpenStudy (anonymous):

what are your answer choices?

OpenStudy (anonymous):

@marcthedark i dont have any. its fill in the blank

OpenStudy (donnie1999):

21 because 9+10 = 21

OpenStudy (anonymous):

damn well i got 1/16

OpenStudy (anonymous):

it would be 1/6 she is only doing it 6 times hahaha

OpenStudy (anonymous):

so that would be a 6.25% chance? @marcthedark

OpenStudy (anonymous):

1/6 hahahah im giving you the answer hahah

OpenStudy (anonymous):

are you sure? @KingCorbin027 i know she is only doing it 6 times, but out of those 6 times,4 are heads? i dont think it would be 1/6

OpenStudy (anonymous):

obviously shes doing it 6 times but thats not the question bro the question is "the probably of getting heads 4 times out of 6 coin tosses and each coin toss has a 50% chance of being heads so explain to me how its 1/6 @KingCorbin027

OpenStudy (anonymous):

ok lets break it down... if you have to flip a coin six times right ok it haves to sides on a coin right heads and tails right ok so that makes it a 50/50 chance that she will land on heads or tails ok so if she wants to get for heads in a row out of six flip 6 is the ?/6 because we now know is there in our answer ok so take 25 from 50 it gives you 25 right ok now you know cant get for heads in a row twice so you go down one just off what you know and it would be 1/6

OpenStudy (anonymous):

@marcthedark @madison.bush

OpenStudy (anonymous):

@ganeshie8 help please

OpenStudy (anonymous):

sorry for the auto corrects

OpenStudy (anonymous):

@KingCorbin027 how if the probability of getting 6 heads in a row is 1/64 bro ? the equation of that is 1/6^6 so the equation for this would make sense if it was 1/6^4

OpenStudy (anonymous):

or just do 4/1*1/64=4/64=1/16

OpenStudy (anonymous):

So thats what you think the answer is? @marcthedark 1/16=6.25%

OpenStudy (anonymous):

that would be my answer, but what class are you doing?

OpenStudy (anonymous):

@marcthedark its algebra 2